a) When 2.00 g of lead is heated with sulfur, 2.30 g of lead (II) sulfide is formed. Calculate the percent composition by mass of each element in lead (II) sulfide.

b)A lab technician find that an 11.5 g of a liquid compound contains 6.00 g of carbon and
1.51 g of hydrogen. The remaining mass is oxygen. Determine the percentage composition
of the compound

c)Calculate the percent composition by mass of all elements in the compound
(NH4 ) 2CO3
1
Expert's answer
2020-06-22T13:34:08-0400

2 grams of lead


2.30 grams of lead (II) sulfide formed


A) lead = (2/2.3)(100)


= 86.95 %


Sulfur = ( 0.30/ 2.30) (100)


= 13.04 %


B) 11.5 grams in a liquid


6 grams of carbon


= (6/11.5)(100)


= 52.17%


1.51grams of hydrogen


=( 1.51/11.5)(100)


= 13.13%


3.99 grams of oxygen


= (3.99/11.5)(100)


= 34.70%

C) mass of ammonium carbonate is equal to 96.09g


(NH4 )2 CO3 = 96.09 grams


N = (28/96.09)(100)

= 29.14%


H = (8/96.09)(100)


= 8.33%


C = (12/96.09)(100)


= 12.49%


O = (48/96.09)(100)


= 49.95%